OpenStudy (anonymous):

x-6y=25 2x-7y=30 how to do this problem using the substitution method?

OpenStudy (anonymous):

x=25+6y now substitute in eq.2

OpenStudy (anonymous):

well i actually did the work but i get x=1 and y=-13/3 but when i plug the number into the equations y does not match

OpenStudy (anonymous):

x=25+6y 2x= 50+12y 50+12y-7y = 30 50+5y=30 5y=-20 y=-4 x= 25 - 24 x=1

OpenStudy (amistre64):

-2(x-6y=25) 2x-7y=30 ----------- 5y = -20 y = -4 x-6y=25 -4 -------- x+24=25 x = 1